AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 17:18:26 2025
Author
*
:
Subject
*
:
<p>Waveforms and discriminator ON. Pulser re-connected to all fees via daisy chain. V=100v, I=8.36uA</p> <p> </p> <p>Pulser peak widths (ch FWHM)</p> <p>aida01 ~ 109</p> <p>aida02 ~ 98</p> <p>aida05 ~ 24</p> <p>aida03 ~ 118</p> <p>aida04 ~ 136</p> <p>Attachments 2, 3 & 4 1.8.W spectra, 20us FSR</p> <p>Attachment 5 FEE64 temps OK</p> <p>Attachment 6 per FEE64 ADC data item rates OK</p> <p>Attachment 7 per FEE64 rate spectra</p> <p> </p> <p>Merger reset following procedure from https://elog.ph.ed.ac.uk/DESPEC/36. Keywords for future elog search NetVar, xfer links</p> <p> </p> <p> </p> <p>Pulser walkthrough R94</p> <p>EG&G Ortec 448 Pulser settings</p> <p>Amplitudes 90,000 - 10,000 @ 10,000 step</p> <p>Attenuators x5 IN</p> <p>t_r 50ns tau_d 50us</p> <p>polarity + (- polarity via Cooknell SA1)</p> <p>Frequency int relay</p> <p>Motors box turned ON, motors pi ON, dlink network connector ON. Motors box, pi and dlink all in 19 inch rack but power supply is from a uk mains extender daisy chained to vacuum mains extender. 19 inch rack power supply remains on clean line.</p> <p>Pulser peak widths (ch FWHM)</p> <p>aida01 ~ 105</p> <p>aida02 ~ 98</p> <p>aida05 ~ 23</p> <p>aida03 ~ 132</p> <p>aida04 ~ 144</p> <p>See attachments 8 & 9</p> <p> </p> <p>Detectors moved in by pnuematic only. Leakage current increased to 9.6 uA (same as leakage current during bias ramping)</p> <p>Pulser Walk R95</p> <p>Pulser settings as above</p> <p>See attachments 10 & 11</p> <p> </p> <p>Detectors moved back out . Motors box, pi and dlink still turned ON. Leakage current to 8.36 uA</p> <p>Pulser Walk R96</p> <p>Pulser settings as above</p> <p>See attachments 1 & 12</p> <p> </p> <p> </p> <p>Jumper removed from nn bias fee adaptor board (aida04). No jumpers on any adaptor boards now. V=100V, I=8.36uA</p> <p>See attachments 14 & 15 - 1.8.W spectra 20us FSR</p> <p>Pulser on 90,000 x5 attenuate</p> <p>aida01 ~ 123, centroid ~14691</p> <p>aida02 ~ 106, centroid ~ 15038</p> <p>aida05 ~ 23, centroid ~ 14414</p> <p>aida03 ~ 125, centroid ~ 50838</p> <p>aida04 ~ 127, centroid ~ 51754</p> <p> </p> <p>Jumper reinstalled on aida04 nn bias. Note only link 1, link2 inaccesible with fee inserted into ERNI. See attachment 13</p> <p>aida01 ~ 123, centroid ~14688</p> <p>aida02 ~ 114, centroid ~ 15034</p> <p>aida05 ~ 23, centroid ~ 14411</p> <p>aida03 ~ 118, centroid ~ 50854</p> <p>aida04 ~ 135, centroid ~ 51697</p> <p> </p> <p>Pulser off. Background run to be left overnight started 14:24 R97. V=100v, I=8.36 uA.</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown